OpenStudy (anonymous):

the height of a ball thrown in the air is found by the equation H=-16Tsquared+48T when the ball hits the ground it has a height of 0. so find how many seconds it takes to hit the ground (solve for T)

OpenStudy (anonymous):

take t common and then the equation will be 16T+48=0

OpenStudy (anonymous):

????

OpenStudy (anonymous):

0=16T^2+48T T(16T+48)=0 16T+48=0 16T=-48 T=-3

OpenStudy (anonymous):

THNX